OpenStudy (anonymous):

The maximum value of y=3sin(1/3theta)-2 is...explanation please

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Consider the maximum of just \(\sin\frac13\theta\) is \(1\). So the maximum of \(3\sin\frac13\theta\) is just \(3\). The \(-2\) then shifts our maximum down to \(1\).